OpenStudy (anonymous):

Simplify. Square root -25 plz help

OpenStudy (anonymous):

5i

OpenStudy (anonymous):

Can u show me how u got that plz?

OpenStudy (anonymous):

\[\sqrt{-25} = \sqrt{25}\sqrt{-1} = 5 \times i\] where i = \[\sqrt{-1}\]

OpenStudy (anonymous):

Ok so the answer is 5i?

OpenStudy (anonymous):

I believe so..

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

basically we're having to use imaginary numbers to describe the negative root, as an equation containing this expression has no real solution. Please see http://en.wikipedia.org/wiki/Complex_number for more information

OpenStudy (anonymous):

\[\sqrt{-25}= \sqrt{25(-1)}= \sqrt{25i^2}= \sqrt{(5i)^2}=5i \]